# Articulate AnyMesh: Open-Vocabulary 3D Articulated Objects Modeling

### v1

@@ Abstract
3D articulated objects modeling has long been a challenging problem, since it requires to capture both accurate surface geometries and semantically meaningful and spatially precise structures, parts, and joints. Existing methods heavily depend on training data from a limited set of handcrafted articulated object categories (e.g., cabinets and drawers), which restricts their ability to model a wide range of articulated objects in an open-vocabulary context. To address these limitations, we propose Articulate AnyMesh, an automated framework that is able to convert any rigid 3D mesh into its articulated counterpart in an open-vocabulary manner. Given a 3D mesh, our framework utilizes advanced Vision-Language Models and visual prompting techniques to extract semantic information, allowing for both the segmentation of object parts and the construction of functional joints. Our experiments show that Articulate AnyMesh can generate large-scale, high-quality 3D articulated objects, including tools, toys, mechanical devices, and vehicles, significantly expanding the coverage of existing 3D articulated object datasets. Additionally, we show that these generated assets can facilitate the acquisition of new articulated object manipulation skills in simulation, which can then be transferred to a real robotic system. Our Github website is https://articulate-anymesh.github.io/.
Artciulated Objects
@@ 1 Introduction
Figure 1: Articulate AnyMesh turns 3D meshes (e.g. retrieved from Objaverse (Deitke et al., 2024)) into articulated objects. Our pipeline is capable of processing a wide range of objects, including daily necessities, furniture, vehicles and even fictional objects.
@@
The gathering of data on a large scale is an emerging research trend in embodied AI and robotics. Large foundation models built for robotics and embodied AI are extremely data-hungry, intensifying the need for large-scale data collection. Compared to collecting data in the real world, gathering data in simulations is significantly faster and more convenient, making it easier to capture various types of ground truth information such as physical states, segmentation masks, depth maps, etc. Existing works have explored many different aspects of how to carry out large-scale data collection in simulations, ranging from asset generation (Chen et al., 2024), scene generation (Yang et al., 2024b, c), task design (Wang et al., 2023), demonstration collection (Dalal et al., 2023; Ha et al., 2023), reward design (Ma et al., 2023), etc.
@@
Despite these efforts, a key challenge remains: collecting diverse and realistic articulated objects essential for everyday life, which is vital for producing diverse data that can be generalized to real-world applications. One intuitive approach to achieve this is to use generative models. While there has been substantial progress in 3D asset generation, few available methods are capable of addressing the demand for the collection of articulated objects. Most 3D generation approaches, utilizing a forward pass of a trained network (Long et al., 2024; Hong et al., 2023; Xu et al., 2024; Shi et al., 2023a, b), or SDS loss optimization (Poole et al., 2022; Wang et al., 2024; Qiu et al., 2024), produce only the surface of the object. These objects can only be manipulated as a whole body. For instance, a closet produced through these 3D generation methods cannot be opened or used to store clothes. Part-aware 3D generation approaches (Gao et al., 2019; Yang et al., 2022; Mo et al., 2019; Wu et al., 2020; Nakayama et al., 2023; Koo et al., 2023; Liu et al., 2024a) generate 3D objects together with their part-specific details. Although these methods are more sensitive to structure, the objects produced are still restricted to whole-object manipulation for the lack of motion parameters. Articulated object creation approaches (Jiang et al., 2022b; Liu et al., 2023a; Chen et al., 2024; Lei et al., 2023; Liu et al., 2024c; Mandi et al., 2024; Nie et al., 2023; Gadre et al., 2021) are capable of producing articulated objects with several interactive parts, demonstrating functionality. However, such methods require dense observation of a to be reconstructed articulated object in multiple joint states (Jiang et al., 2022b; Liu et al., 2023a; Mandi et al., 2024; Huang et al., 2021), or are restricted to the limited-scale data and object categories used to train their network (Chen et al., 2024; Lei et al., 2023; Liu et al., 2024c). As a result, current methods struggle to automatically generate a wide variety of articulated objects, particularly those from underrepresented or absent categories in existing datasets (Xiang et al., 2020; Wang et al., 2019; Geng et al., 2023).
@@
Consequently, the collection of diverse articulated objects presents extra challenges compared to non-articulated 3D assets: (1) accurate semantic structures must accompany geometry and appearance, (2) articulation parameters such as joint orientation and position are required, and (3) the relatively small-scale articulated object datasets, compared to 3D object datasets, are insufficient to support the training of a generalizable model. To overcome these challenges, we introduce Articulate AnyMesh, an automated pipeline that converts any 3D mesh into a corresponding articulated asset. In order to go beyond existing articulated object datasets, our pipeline leverages prior knowledge from visual and language foundation models (Achiam et al., 2023; Kirillov et al., 2023; Rombach et al., 2022) and generalizable geometric clues, rather than relying on existing labeled articulated object data. Our pipeline starts with a 3D mesh, either generated or handcrafted, followed by the stages of Movable Part Segmentation, Articulation Estimation and Refinement.
@@
In Movable Part Segmentation, the goal is to identify all movable parts and determine their semantics for subsequent articulation estimation. Recent open-vocabulary 3D part segmentation methods (Liu et al., 2023b; Zhou et al., 2023; Umam et al., 2024; Xue et al., 2023; Yang et al., 2024a; Liu et al., 2024a) utilize an open-vocabulary segmentation model (namely SAM) and a multi-modal foundation model on rendered 2D images, integrating the 2D information to achieve 3D part segmentation. In this work, we adopt the pipeline proposed by PartSlip++ (Zhou et al., 2023) for 3D part segmentation.
@@
After obtaining the 3D segmentation of non-fixed parts and their semantics, the next stage in our proposed pipeline focuses Articulation Estimation. Unlike previous approaches (Li et al., 2020; Huang et al., 2021; Zeng et al., 2024), which rely on training datasets to predict articulation parameters. To overcome the limitations of existing datasets and extent to open-vocabulary manner, we extract geometric clues inherent in articulated objects instead of relying on learned models. Specifically, the joint connecting two segmented parts is inherently related to the geometry of the connected area. For example, in cases where the connection follows a straight line, such as a laptop hinge or an open door, this line effectively represents the joint since the hinge structures align with it. We provide GPT-4o with such information through visual prompting, allowing it to leverage common-sense knowledge geometric reasoning to infer joint parameters. This approach allows generalization to unseen objects with accurate segmentation, and the performance will continue to improve as the vision-language model becomes more advanced.
@@
At this stage, a functional articulated object has already been created. However, if the input mesh is a generated or scanned surface mesh that only contains surface geometry and texture, the segmented 3D parts often suffer from occlusion, and may leave holes in other parts. For example, in the closed state, only the outer surface of the drawer will be segmented, and its internal structure will not be present at all.
To address these issues and complete our proposed pipeline, we incorporated a final stage: Refinement. In this stage, we leverage 2D diffusion models (Rombach et al., 2022; Poole et al., 2022) to close holes, enhance geometry, and improve texture quality. Such models are trained on millions of images, establishing a solid prior about 3D objects. This 3D prior can be effectively distilled using SDS (Score Distillation Sampling) loss (Poole et al., 2022). Although this optimization process is not aware of 3D part segmentation, causing parts to grow into each other, we solve this by a simple yet effective optimization strategy that incorporates random transformations of movable parts according to their joint parameters during the SDS loss optimization phase.
@@
We conducted experiments to compare joint parameter estimation accuracy and unconditional generation capability with various baselines. Experimental results demonstrate that while our method matches state-of-the-art articulated object modeling methods for selected categories that they are trained on within PartNet-Mobility dataset (Xiang et al., 2020), it also exhibits the capacity to model a broader range of articulated objects beyond these categories. We summarize our contributions as follows:
@@
  • We propose a novel optimization strategy that enables using diffusion prior to refine articulated objects while keeping the semantics of parts from being changed or lost.
@@ 2.1 Articulated object modeling
One line of work focuses on the perception end, including reconstruction of articulated objects from observations, joint parameter estimation, etc. (Jiang et al., 2022b) and (Huang et al., 2021) train a network to reconstruct an articulated object from input multi-frame point clouds. Some other works (Hsu et al., 2023; Gadre et al., 2021; Nie et al., 2023; Wang et al., 2022) learn joint parameters from active interaction. (Liu et al., 2023a), (Weng et al., 2024), (Wei et al., 2022) and (Mandi et al., 2024) reconstruct the geometry and joint parameters of an articulated object from multiview images obtained at distinct states of the articulated object. (Chen et al., 2024) reconstructs articulated objects from a single real-world RGB image through a network trained on large-scale synthetic data. (Li et al., 2020; Jiang et al., 2022a; Sun et al., 2024; Yan et al., 2020; Liu et al., 2022, 2023c; Geng et al., 2023) estimates joint parameters and part segments given the poincloud or image of an articulated object. However, these approaches either demand extensive observation of the same object or rely on existing articulated object data to train neural networks, with limited ability to generalize across only a small range of categories.
Alongside perception works, generation works including (Lei et al., 2023) and (Liu et al., 2024c), represents articulated objects as graphs and employs a diffusion model to fit the distribution. The shape, position, and semantic information of the parts are encoded in the vertices, while joint parameters are stored in the edges. Similar to perception methods, these generation methods also depend on existing articulated object datasets to train their diffusion models and are unable to generalize beyond the training data.
Additionally, there is a recent survey on articulated object modeling (Liu et al., 2024b).
@@ 2.2 Part aware 3D generation
Instead of generating an 3D object as a whole, part aware 3D generation methods generate objects with part-level information. (Gao et al., 2019) learns two separate VAEs to model global structural information and detailed part geometries. (Yang et al., 2022) encodes 3D objects into a representation that disentangles structure and geometry. (Mo et al., 2019) utilizes a graph VAE to encode shapes into hierarchical graph representations. (Wu et al., 2020) applies a Seq2Seq generative network for part assembly and generation. (Nakayama et al., 2023) independently models part geometry and global part transformation, and conditioned on both of them, a diffusion model synthesizes the final shape. (Koo et al., 2023), on the other hand, generates part geometry latents with one diffusion model and synthesizes the global geometry with another diffusion model conditioned on these latents. (Liu et al., 2024a), a recent release, learns a Neus model from generated multi-view images and corresponding 2D segmentation maps provided by (Kirillov et al., 2023), and extracts 3D segmentation masks using a clustering algorithm.
@@ 3 Method
Figure 2: Articulate AnyMesh converts 3D meshes from various sources to high quality articulated objects via three main parts: A. Movable Part Segmentation, B. Articulation Estimation, C. Refinement.
@@ 3.1 Overview
Our pipeline converts any mesh into its articulated counterpart, including hand-crafted meshes with part geometry, surface meshes generated through text-to-3D or image-to-3D methods and meshes reconstructed from real-world objects. The process involves three main steps, including Movable Part Segmentation, Articulation Estimation and Refinement, as depicted in Figure 2. Specifically, the Movable Part Segmentation step employs the method proposed by PartSlip++ (Liu et al., 2023b) to segment movable parts from the input mesh. Articulation Estimation extracts the articulation parameters for each movable part by leveraging geometric cues and prior knowledge from visual and language foundation models. In the Refinement stage, the geometry and texture of the articulated object are enhanced with geometry-aware 2D diffusion model Richdreamer (Qiu et al., 2024), guided by SDS loss to distill the diffusion model’s 3D prior. During this stage, movable parts undergo random transformation to improve optimization performance.
@@ 3.2 Movable Part Segmentation via VLM assistant
As shown in Figure 2A, this step aims to segment all movable parts from a 3D mesh and identify their semantics in an open-vocabulary manner. We primarily use PartSlip++ (Zhou et al., 2023) as our segmentation method. In PartSlip++, an image of the input mesh is first fed into a VLM to recognize all movable parts. Next, an open-vocabulary grounding model (Ren et al., 2024), and SAM are applied to 2D rendered images of the input mesh to extract bounding boxes and segmentation masks of the relevant part semantic labels. These multiview 2D detection and segmentation results are then fused to produce the 3D part segmentation.
@@ 3.3 Articulation Estimation via Geometry-aware Visual Prompting
In some cases, a hinge may not be positioned on the object’s surface, resulting in only one end of the rotation axis identifiable on the object’s surface, such as with knobs. In these cases, we assume the rotation axis to be perpendicular to the plane fitted to the connected area. The axis is then determined by the normal vector of this plane and the single selected point. When the hinge mechanism is hidden beneath the surface, the connecting structure cannot be assumed to be entirely within the connecting area. Thus, in addition to candidate points from the connecting area, we sample additional points from the part surface. Specifically, we use the $l_{0}$-cut pursuit (Landrieu & Obozinski, 2017) algorithm to generate super points from the part’s point cloud, using point normals and color as features, and derive candidate points from these super points. This ensures the new candidate points are distributed across different geometric and textured regions of the part.
@@
3.4 Refinement via Randomized Joint State
After completing the stages of 3D segmentation and joint estimation, we obtain an articulated object with the correct structure. However, when the input mesh is a surface mesh, the result often remains flawed. Object parts segmented from surface meshes—lacking inner structures—inevitably suffer from occlusion, leading to incomplete part meshes with holes.
To address this, we integrate a refinement step into our pipeline. This step completes occluded geometries, fills holes, and enhances texture quality by distilling 3D object geometry and texture priors from a geometry-aware 2D diffusion model into the articulated object using SDS loss (Poole et al., 2022). The refinement process builds on Richdreamer (Qiu et al., 2024), which sequentially applies a normal-depth diffusion model and an albedo diffusion model to optimize the geometry and texture of 3D object, respectively.
However, directly applying existing SDS optimization method (Qiu et al., 2024) on articulated objects in fixed postures often causes parts to grow and overlap, resulting in artifacts when the object parts move. To solve this problem, we propose a novel optimization technique for articulated objects: randomizing the articulation configuration during optimization. In each optimization step, along with randomly sampling camera positions and diffusion timesteps, we also sample motion parameters (e.g., rotation angles and translation lengths) for each non-fixed part and apply these transformations. This approach ensures that most parts are exposed in certain poses, while overlapping geometry in one pose becomes evident as artifacts in other poses, allowing gradual optimization. Since the parts are already recognizable at initialization, the diffusion model preserves their semantics effectively, ensuring that the meaning and identity of the parts remain intact throughout the refinement process.
For implementation details please refer to appendix A.

### v2

@@ Abstract
3D articulated objects modeling has long been a challenging problem, since it requires to capture both accurate surface geometries and semantically meaningful and spatially precise structures, parts, and joints. Existing methods heavily depend on training data from a limited set of handcrafted articulated object categories (e.g., cabinets and drawers), which restricts their ability to model a wide range of articulated objects in an open-vocabulary context. To address these limitations, we propose Articulate AnyMesh, an automated framework that is able to convert any rigid 3D mesh into its articulated counterpart in an open-vocabulary manner. Given a 3D mesh, our framework utilizes advanced Vision-Language Models and visual prompting techniques to extract semantic information, allowing for both the segmentation of object parts and the construction of functional joints. Our experiments show that Articulate AnyMesh can generate large-scale, high-quality 3D articulated objects, including tools, toys, mechanical devices, and vehicles, significantly expanding the coverage of existing 3D articulated object datasets. Additionally, we show that these generated assets can facilitate the acquisition of new articulated object manipulation skills in simulation, which can then be transferred to a real robotic system. Our Github website is https://articulateanymesh.github.io/.
@@ 1 Introduction
Figure 1: Articulate AnyMesh turns 3D meshes (e.g. retrieved from Objaverse [1]) into articulated objects. Our pipeline is capable of processing a wide range of objects, including daily necessities, furniture, vehicles and even fictional objects.
@@
The gathering of data on a large scale is an emerging research trend in embodied AI and robotics. Large foundation models built for robotics and embodied AI are extremely data-hungry, intensifying the need for large-scale data collection. Compared to collecting data in the real world, gathering data in simulations is significantly faster and more convenient, making it easier to capture various types of ground truth information such as physical states, segmentation masks, depth maps, etc. Existing works have explored many different aspects of how to carry out large-scale data collection in simulations, ranging from asset generation [2], scene generation [3, 4, 5], task design [6], demonstration collection [7, 8], reward design [9], etc.
@@
Despite these efforts, a key challenge remains: collecting diverse and realistic articulated objects essential for everyday life, which is vital for producing diverse data that can be generalized to real-world applications. One intuitive approach to achieve this is to use generative models. While there has been substantial progress in 3D asset generation, few available methods are capable of addressing the demand for the collection of articulated objects. Most 3D generation approaches, utilizing a forward pass of a trained network [10, 11, 12, 13, 14], or SDS loss optimization [15, 16, 17], produce only the surface of the object. These objects can only be manipulated as a whole body. For instance, a closet produced through these 3D generation methods cannot be opened or used to store clothes. Part-aware 3D generation approaches [18, 19, 20, 21, 22, 23, 24] generate 3D objects together with their part-specific details. Although these methods are more sensitive to structure, the objects produced are still restricted to whole-object manipulation for the lack of motion parameters. Articulated object creation approaches [25, 26, 2, 27, 28, 29, 30, 31] are capable of producing articulated objects with several interactive parts, demonstrating functionality. However, such methods require dense observation of a to be reconstructed articulated object in multiple joint states [25, 26, 29, 32], or are restricted to the limited-scale data and object categories used to train their network [2, 27, 28]. As a result, current methods struggle to automatically generate a wide variety of articulated objects, particularly those from underrepresented or absent categories in existing datasets [33, 34, 35].
@@
Consequently, the collection of diverse articulated objects presents extra challenges compared to non-articulated 3D assets: (1) accurate semantic structures must accompany geometry and appearance, (2) articulation parameters such as joint orientation and position are required, and (3) the relatively small-scale articulated object datasets, compared to 3D object datasets, are insufficient to support the training of a generalizable model. To overcome these challenges, we introduce Articulate AnyMesh, an automated pipeline that converts any 3D mesh into a corresponding articulated asset. In order to go beyond existing articulated object datasets, our pipeline leverages prior knowledge from visual and language foundation models [36, 37, 38] and generalizable geometric clues, rather than relying on existing labeled articulated object data. Our pipeline starts with a 3D mesh, either generated or handcrafted, followed by the stages of Movable Part Segmentation, Articulation Estimation and Post-Processing.
@@
In Movable Part Segmentation, the goal is to identify all movable parts and determine their semantics for subsequent articulation estimation. Recent open-vocabulary 3D part segmentation methods [39, 40, 41, 42, 43, 24] utilize an open-vocabulary segmentation model (namely SAM) and a multi-modal foundation model on rendered 2D images, integrating the 2D information to achieve 3D part segmentation. In this work, we adopt the pipeline proposed by PartSlip++ [40] for 3D part segmentation.
@@
After obtaining the 3D segmentation of non-fixed parts and their semantics, the next stage in our proposed pipeline focuses Articulation Estimation. Unlike previous approaches [44, 32, 45], which rely on training datasets to predict articulation parameters. To overcome the limitations of existing datasets and extent to open-vocabulary manner, we extract geometric clues inherent in articulated objects instead of relying on learned models. Specifically, the joint connecting two segmented parts is inherently related to the geometry of the connected area. For example, in cases where the connection follows a straight line, such as a laptop hinge or an open door, this line effectively represents the joint since the hinge structures align with it. We provide GPT-4o with such information through visual prompting, allowing it to leverage common-sense knowledge geometric reasoning to infer joint parameters. This approach allows generalization to unseen objects with accurate segmentation, and the performance will continue to improve as the vision-language model becomes more advanced.
@@
At this stage, a functional articulated object has been created. However, if the input mesh is a generated or scanned surface mesh that contains only surface geometry and texture, the segmented 3D parts often suffer from occlusion and may exhibit holes. For example, in the closed state, only the outer surface of a microwave may be segmented, with the geometry of its internal cavity entirely missing. To address these issues and complete the pipeline, we incorporate an optional final stage: Post-Processing. Here, we leverage 3D generative models [46, 47] to perform shape completion (filling holes and recovering missing geometry) and texture generation, making the pipeline self-contained for surface meshes.
@@
We conducted experiments to compare joint parameter estimation accuracy with various baselines. Experimental results demonstrate that while our method matches state-of-the-art articulated object modeling methods for selected categories that they are trained on within PartNet-Mobility dataset [33], it also exhibits the capacity to model a broader range of articulated objects beyond these categories. We summarize our contributions as follows:
@@
  • We conducted extensive experiments to evaluate the performance of our pipeline. We compared our method with others in joint parameter estimation and outperformed previous approaches in both in-domain and out-of-domain settings. We also carried out manipulation policy learning experiments to demonstrate the usefulness of the data collected by our pipeline.
@@ 2.1 Articulated object modeling
One line of work focuses on the perception end, including reconstruction of articulated objects from observations, joint parameter estimation, etc. [25] and [32] train a network to reconstruct an articulated object from input multi-frame point clouds. Some other works [48, 31, 30, 49] learn joint parameters from active interaction. [26], [50], [51] and [29] reconstruct the geometry and joint parameters of an articulated object from multiview images obtained at distinct states of the articulated object. [2] reconstructs articulated objects from a single real-world RGB image through a network trained on large-scale synthetic data. [44, 52, 53, 54, 55, 56, 35] estimates joint parameters and part segments given the poincloud or image of an articulated object. However, these approaches either demand extensive observation of the same object or rely on existing articulated object data to train neural networks, with limited ability to generalize across only a small range of categories.
Alongside perception works, generation works including [27], [28] and [57], represents articulated objects as graphs and employs a diffusion model to fit the distribution. The shape, position, and semantic information of the parts are encoded in the vertices, while joint parameters are stored in the edges. Similar to perception methods, these generation methods also depend on existing articulated object datasets to train their diffusion models and are unable to generalize beyond the training data.
@@ 2.2 Part aware 3D generation
Instead of generating an 3D object as a whole, part aware 3D generation methods generate objects with part-level information. [18] learns two separate VAEs to model global structural information and detailed part geometries. [19] encodes 3D objects into a representation that disentangles structure and geometry. [20] utilizes a graph VAE to encode shapes into hierarchical graph representations. [21] applies a Seq2Seq generative network for part assembly and generation. [22] independently models part geometry and global part transformation, and conditioned on both of them, a diffusion model synthesizes the final shape. [23], on the other hand, generates part geometry latents with one diffusion model and synthesizes the global geometry with another diffusion model conditioned on these latents. [24] learns a Neus model from generated multi-view images and corresponding 2D segmentation maps provided by [37], and extracts 3D segmentation masks using a clustering algorithm. [59] starts from multi-view images, applies diffusion-based networks for consistent part segmentation, completes each part across views, and reconstructs 3D parts using a pre-trained model..
@@ 3 Method
Figure 2: Articulate AnyMesh converts 3D meshes from various sources to high quality articulated objects via three main parts: A. Movable Part Segmentation, B. Articulation Estimation, C. Post-Processing.
@@ 3.1 Overview
Our pipeline converts any mesh into its articulated counterpart, including hand-crafted meshes with part geometry, surface meshes generated through text-to-3D or image-to-3D methods and meshes reconstructed from real-world objects. The process involves three main steps, including Movable Part Segmentation, Articulation Estimation and Post-Processing, as depicted in Figure 2. Specifically, the Movable Part Segmentation step employs the method proposed by PartSlip++ [39] to segment movable parts from the input mesh. Articulation Estimation extracts the articulation parameters for each movable part by leveraging geometric cues and prior knowledge from visual and language foundation models. The optional Post-Processing stage enhances geometry and texture using an off-the-shelf 3D shape completion model [46] and an off-the-shelf texture generation model [47].
@@ 3.2 Movable Part Segmentation via VLM assistant
As shown in Figure 2A, this step segments all movable parts from a 3D mesh and identifies their semantics in an open-vocabulary manner. We primarily adopt PartSlip++ [40] for segmentation. Given a 3D object and a list of part labels, PartSlip++ applies an open-vocabulary grounding model [60, 61] and SAM to 2D rendered images to extract bounding boxes and segmentation masks for the relevant semantic labels. These multi-view 2D detections are then fused to produce the final 3D part segmentation. In our pipeline, the label list is generated by feeding an image of the input mesh into a VLM to identify all movable parts.
@@ 3.3 Articulation Estimation via Geometry-aware Visual Prompting
In some cases, a hinge may not be positioned on the object’s surface, resulting in only one end of the rotation axis identifiable on the object’s surface, such as with knobs. In these cases, we assume the rotation axis to be perpendicular to the plane fitted to the connected area. The axis is then determined by the normal vector of this plane and the single selected point. When the hinge mechanism is hidden beneath the surface, the connecting structure cannot be assumed to be entirely within the connecting area. Thus, in addition to candidate points from the connecting area, we sample additional points from the part surface. Specifically, we use the $l_{0}$-cut pursuit [62] algorithm to generate super points from the part’s point cloud, using point normals and color as features, and derive candidate points from these super points. This ensures the new candidate points are distributed across different geometric and textured regions of the part.
@@
3.4 Geometry and Texture Post-Processing
After completing 3D segmentation and joint estimation, we obtain an articulated object with the correct articulation structure. However, when the input is a surface mesh, the result often remains incomplete. Parts segmented from surface meshes—lacking internal geometry—are prone to occlusions, resulting in incomplete meshes with holes. To address this, we introduce an optional post-processing step for surface meshes (primarily generated ones), where off-the-shelf 3D generation models are used for shape completion and texture generation, as illustrated in Section C of Figure 2.
For shape completion, we use HoloPart[46], which takes a mesh with 3D segmentation labels as input and outputs mesh parts with completed geometry. As shown in the pipeline figure, HoloPart repairs issues such as holes in the bucket base (caused by occlusion from the handle) and a broken handle. For texture generation, we use Meshy[47] to apply textures to the completed parts. As shown in the figure, Meshy can generate high-quality textures with diverse materials, such as wood, plastic, and metal.